how much must you deposit in an account that pays 6.25% interest, compounded annually, to have a balance of 700 after two years answer?
Answer by Cromlix(4381)

Hi there,
100% + 6.25% = 106.25%
106.25% = 106.25/100 = 1.0625
x * 1.0625^2 = 700 (* = times)
x = 700/1.0625^2
x = 620.07 (2 decimal figures)
Hope this helps :-)